AD Use the method of Laplace transforms and the accompanying proof results to solve the initial value problem. y + 6y + 8y
Answer #1

G o 2 satco 10 1-elos [ (g(x)) = švézt geht = (1-82t)? Solution: Given differential equation is y +6y! +8y = f(t); 7(0) =0,